# SATURN: SAT-based Reinforcement Learning to Unleash Language Model Reasoning

### v1

@@ Abstract
We introduce SATURN-2.6k, a dataset of 2,660 SAT problems with varying difficulty. It supports the evaluation of how LLM reasoning changes with problem difficulty. We apply SATURN to DeepSeek-R1-Distill-Qwen and obtain SATURN-1.5B and SATURN-7B. We achieve several notable results: ❶ On SAT problems, SATURN-1.5B and SATURN-7B achieve average pass@3 improvements of +14.0 and +28.1, respectively. ❷ On math and programming tasks, SATURN-1.5B and SATURN-7B improve average scores by +4.9 and +1.8 on benchmarks (e.g., AIME, LiveCodeBench). ❸ Compared to the state-of-the-art (SOTA) approach in constructing RL tasks, SATURN achieves further improvements of +8.8%. We release the source code, data, and models to support future research [4].
@@ 1 Introduction
Recently, reinforcement learning (RL) has become a promising paradigm for unleashing the reasoning capability of large language models (LLMs), particularly in math, programming, and logical reasoning (e.g., OpenAI-o1 [20], DeepSeek-R1 [12], Kimi-k1.5 [35]). During the RL training process, the design of RL tasks plays a critical role [15, 31, 30]. A well-designed RL task should elicit LLMs’ reasoning capability, fostering behaviors such as hesitation, reflection, backtracking, summarization, and verification [41, 42, 51, 30, 27].
However, how to design RL tasks that can continuously enhance LLMs’ reasoning capability remains an open question. We think a well-designed RL task for reasoning should satisfy the following three criteria: ❶ Scalability. RL training requires large-scale data. RL tasks should support scalable data without human annotation or expensive LLMs’ synthesis. ❷ Verifiability. RL rewards must be unambiguously correct. The outputs of LLMs for the task should be easy to verify. ❸ Controllable Difficulty. Reasoning capability emerges progressively [40]. RL tasks should support the difficulty control to enable curriculum learning, allowing LLMs to gradually develop more complex reasoning skills [16].
@@
ScaleQuest [13]
GSM8K (Math) [10]
LiveCodeBench [21]
Game Werewolf [43, 46]
LMRL Gym [6]
SPAG [9]
Knights&Knaves [44]
@@
Table 1 shows the features of current mainstream RL tasks. None of them satisfy all three criteria. Existing RL tasks can be divided into two categories: (1) One category of RL tasks requires LLMs to solve math or programming problems, with rewards based on the correctness of the final answer or code [8, 26, 7, 10]. However, these tasks rely on human annotation for ground-truth solutions or test cases, suffer from a lack of high-quality problems, and offer only coarse control over reasoning difficulty [22, 25]. (2) Another category focuses on manually designed reasoning tasks [13, 43, 46, 44]. For instance, Logic-RL [45] leverages natural language logic K&K puzzles to improve LLMs’ reasoning capability through RL. However, they also present limitations, such as hard to scale up due to reliance on sampling from LLMs [9, 13], hard to verify despite relying on LLMs for cross-validation [49, 13, 39], and hard to control difficulty [43, 46].
@@
To this end, we propose Boolean Satisfiability (SAT) problem as the task for RL. Figure 1 shows an illustration of SAT problems and corresponding features. SAT satisfies all three desiderata we outlined earlier: ❶Scalability. SAT instances can be generated programmatically at scale without human annotation or LLM synthesis, allowing for virtually unlimited training data. ❷ Verifiability. SAT is a well-established NP-complete problem in theoretical computer science [11]. The correctness of a solution can be easily verified in linear time. But solving SAT problems requires complex reasoning. ❸ Controllable Difficulty. The difficulty of SAT instances can be precisely adjusted (e.g., number of variables, clauses), making it suitable for curriculum learning. What’s more, SAT serves as a universal substrate for limited forms of logical reasoning, as many problems in propositional logic, finite-domain first-order logic, and modal logic can be systematically reduced to SAT [17, 34, 23].
@@
We apply SATURN to DeepSeek-R1-Distill-Qwen-1.5B and 7B [12], obtaining SATURN-1.5B and SATURN-7B. Experiments show that SATURN effectively enhances LLMs’ reasoning capability in generalizable scenarios:
@@
  • The reasoning capability learned from SATURN transfers well to math and programming tasks, bringing average improvements of +4.9 and +1.8 on popular benchmarks such as AIME [3], AMC [2], MATH-500 [19], GPQA Diamond [33], and LiveCodebench [21] for two SATURN models.
@@ 2.1 SATURN Learning Loop Framework
[…]
@@ 2.2 SAT Instances Construction
In this subsection, we formalize the construction of SAT instances. A SAT problem determines if a propositional formula can be satisfied by a Boolean truth assignment. Formally, we define a $(n,k,l)$-SAT instance in conjunctive normal form (CNF) as:
$\Phi=\ &\left(x_{a_{1,1}}\lor\neg x_{a_{1,2}}\lor\dots\lor x_{a_{1,n}}\right)\land\cdots\land\left(x_{a_{l,1}}\lor\dots\lor\neg x_{a_{l,n}}\right)\\ \text{where }&a_{i,j}\in\{1,\dots,k\},\quad i\in[1,l]\cap\mathbb{N},\ j\in[1,n]\cap\mathbb{N}\end{aligned}\right.$ (1)
where each clause contains exactly $n$ variables (literals), each being either $x_{i}$ or its negation $\neg x_{i}$, $k$ is the total number of variables, and $l$ is the total number of clauses. Based on the definition, we design a SAT instance constructor, SAT_Construction$(n,k,l,m)$, which uniformly samples $m$ SAT instances from the space of $(n,k,l)$-SAT. By adjusting the parameters $(n,k,l,m)$, SAT_Construction enables scalable and controllable construction of SAT instances. The design details of the constructor algorithm are provided in Appendix B.
@@ 2.3 Estimation of Task Difficulty
As a canonical NP-complete problem [11], SAT admits a polynomial-time reduction from any other NP problem [24]. SAT exhibits a known phase transition phenomenon: when the clause-to-variable ratio $\alpha_{c}=l/k$ approaches a critical threshold (typically near $4.26$ for 3-SAT), the probability of satisfiability drops sharply, and problem difficulty peaks. This phenomenon probably stems from replica symmetry breaking (RSB) [52]: near $\alpha_{c}$, the solution space fractures into disconnected clusters separated by energy barriers. Beyond $\alpha_{c}$, the space collapses, reducing complexity.
However, RSB theory is designed for heuristic SAT solvers. For humans or LLMs solving SAT problems through logical steps like trial, verification, and reasoning, such solver-like phase transitions are hardly observable in human-like thinking processes. While any $n$-SAT ($n>3$) can be reduced to 3-SAT [24], they differ significantly for LLMs in solution space size and token length.
Prior work [18] on SAT tasks for LLMs typically categorized difficulty based on phase transition points. To systematically estimate task difficulty, we define an analytical estimator of the expected solution space size. Given a $(n,k,l)$ SAT instance, its difficulty for LLMs can be approximately estimated by:
@@
Eq. (2) provides a more controllable, fine-grained estimation of SAT task difficulty. The detailed derivation is provided in Appendix C. To further validate Eq. (2), we evaluate LLMs’ performance on SAT instances with varying difficulty levels. As shown in Figure 3, each point represents a LLM’s average pass@3 on the same estimated difficulty instances. Pass@3 generally decreases as $D(n,k,l)$ increases, suggesting that our estimation aligns with the solvability trends observed in practical LLMs. Stronger LLMs maintain higher pass@3, while weaker LLMs exhibit lower scores overall.
@@ 2.4 Reinforcement Learning with GRPO
In this subsection, we introduce the single-stage RL training for given $(n,k,l)$-difficulty tasks. RL can further improve LLMs’ generalization by directly optimizing policy gradients over diverse reasoning trajectories [12]. Given the SAT tasks, we then train LLMs using the original sample-level GRPO to optimize the policy $\pi_{\theta}$ with KL divergence penalty. The GRPO objective function is defined as:
@@
where $\text{where }r_{i,t}(\theta)$ $=\frac{\pi_{\theta}(o_{i,t}\mid q,o_{i,<t})}{\pi_{\theta_{\text{old}}}(o_{i,t}\mid q,o_{i,<t})},\quad\hat{A}_{i,t}=\frac{r_{i}-\text{mean}(\mathbf{r})}{\text{std}(\mathbf{r})}$
@@ 3.2 Experimental Setup
❶ Building upon SAT_Construction tool and difficulty estimation, we release SATURN-2.6k, a curated benchmark designed to evaluate LLMs’ reasoning capability across varying complexity. SATURN-2.6k consists of 1,500 training instances and 160 test instances that share the same estimated difficulty level. To assess performance under increasing task complexity, SATURN-2.6k further includes 1,000 test instances from 10 unseen harder difficulty levels, with 100 instances per level. These levels are selected based on our difficulty estimation $D(n,k,l)$, enabling a systematic analysis of how LLM performance changes as problem difficulty increases. Additionally, custom datasets of desired difficulty can be constructed using our open-sourced SAT_Construction tool. ❷ For math and programming tasks, following DeepSeek-AI [12], we use AIME 24/25 [3], AMC 22/23 [2], MATH-500 [19], GPQA Diamond [33], and LiveCodeBench v4_v5 subset [21].
@@
We conduct evaluations against several 1.5B and 7B parameter reasoning models as the baselines, which include DeepSeek-R1-Distill-Qwen-1.5B & 7B [12], Still-3-1.5B-Preview [38], s1.1-1.5B & 7B [30], z1-7B [48], OpenThinker-7B [36], and DeepScaleR-1.5B-Preview [28].
@@
[…]
@@ 3.3 RQ1: SATURN Substantially Improves Performance on SAT Tasks
We evaluate the performance of SATURN-1.5B and SATURN-7B on SAT tasks using SATURN-2.6k test set. Specifically, the evaluation involves unseen SAT instances that were not included in the training data. The results, presented in Table 2, 812, and detailed in Appendix G, demonstrate the performance of LLMs across different SAT difficulties.
@@ 3.4 RQ2: SATURN Demonstrates Strong Generalization to Math and Programming
SATURN shows strong generalization to math and programming tasks. On the AIME 24/25 benchmark, SATURN-1.5B outperforms z1-7B by 8.3 and s1.1-7B by 21.7. Similarly, SATURN-7B achieves a strong improvement on the Math500 dataset, increasing from 93.2 to 95.0. On LiveCodeBench, it improves from 35.4 to 37.7. On average, SATURN-1.5B improves by +4.9, and SATURN-7B improves by +1.8 across these benchmarks. These results highlight that SATURN enhances the reasoning performance of LLMs across various math and programming tasks, demonstrating strong generalization of the learned reasoning capabilities from SAT.
@@ 3.5 RQ3: SATURN Serves as a Complement and Further Enhances LLM Reasoning
RQ3 studies the relationship between SATURN and existing RL tasks. Beyond the DeepSeek-R1-Distill-Qwen-7B, we introduce two additional models: Qwen2.5-7B-Instruct-1M [47, 37] following Logic-RL [45] settings, and DeepScaleR-1.5B-Preview [28], which is further RL trained on 40k math and programming examples from DeepSeek-R1-Distill-Qwen-7B. We compare SATURN against several prior constructing RL task approaches, including Logic-RL [45], SPGA [9], and ScaleQuest [13], which represent strong baselines. Each approach is applied to different models for comparison. Results are summarized in Table 4.

### v3

@@ Abstract
We introduce SATURN-2.6k, a dataset of 2,660 SAT problems with varying difficulty. It supports the evaluation of how LLM reasoning changes with problem difficulty. We apply SATURN to DeepSeek-R1-Distill-Qwen and obtain SATURN-1.5B and SATURN-7B. We achieve several notable results: ❶ On SAT problems, SATURN-1.5B and SATURN-7B achieve average pass@3 improvements of +14.0 and +28.1, respectively. ❷ On math and programming tasks, SATURN-1.5B and SATURN-7B improve average scores by +4.9 and +1.8 on benchmarks (e.g., AIME, LiveCodeBench). ❸ Compared to the state-of-the-art (SOTA) approach in constructing RL tasks, SATURN achieves further improvements of +8.8%. We release the source code, data, and models to support future research at https://github.com/gtxygyzb/Saturn-code.
@@ 1 Introduction
Recently, reinforcement learning (RL) has become a promising paradigm for unleashing the reasoning capability of large language models (LLMs), particularly in math, programming, and logical reasoning (e.g., OpenAI-o1 [openai_o1], DeepSeek-R1 [deepseek_r1], Kimi-k1.5 [kimi_k1.5]). During the RL training process, the design of RL tasks plays a critical role [openai_o3, openai_o3_web, s1]. A well-designed RL task should elicit LLMs’ reasoning capability, fostering behaviors such as hesitation, reflection, backtracking, summarization, and verification [cot, self_verification, stepback, s1, Verify_Step].
However, how to design RL tasks that can continuously enhance LLMs’ reasoning capability remains an open question. We think a well-designed RL task for reasoning should satisfy the following three criteria: ❶ Scalability. RL training requires large-scale data. RL tasks should support scalable data without human annotation or expensive LLMs’ synthesis. ❷ Verifiability. RL rewards must be unambiguously correct. The outputs of LLMs for the task should be easy to verify. ❸ Controllable Difficulty. Reasoning capability emerges progressively [Curriculum_Learning]. RL tasks should support the difficulty control to enable curriculum learning, allowing LLMs to gradually develop more complex reasoning skills [difficulty_controllable].
@@
ScaleQuest [question_synthesis]
GSM8K (Math) [gsm8k]
LiveCodeBench [lcb]
Game Werewolf [wolf1, wolf2]
LMRL Gym [lmrl]
SPAG [Self-playing-Adversarial]
Knights&Knaves [kk2]
@@
Table 1 shows the features of current mainstream RL tasks. None of them satisfy all three criteria. Existing RL tasks can be divided into two categories: (1) One category of RL tasks requires LLMs to solve math or programming problems, with rewards based on the correctness of the final answer or code [humaneval, aplhacode, mbpp, gsm8k]. However, these tasks rely on human annotation for ground-truth solutions or test cases, suffer from a lack of high-quality problems, and offer only coarse control over reasoning difficulty [aixcoder, aixcoderv2]. (2) Another category focuses on manually designed reasoning tasks [question_synthesis, wolf1, wolf2, kk2]. For instance, Logic-RL [Logic-RL] leverages natural language logic K&K puzzles to improve LLMs’ reasoning capability through RL. However, they also present limitations, such as hard to scale up due to reliance on sampling from LLMs [Self-playing-Adversarial, question_synthesis], hard to verify despite relying on LLMs for cross-validation [codedpo, question_synthesis, puzzbench], and hard to control difficulty [wolf1, wolf2].
@@
To this end, we propose Boolean Satisfiability (SAT) problem as the task for RL. Figure 1 shows an illustration of SAT problems and corresponding features. SAT satisfies all three desiderata we outlined earlier: ❶Scalability. SAT instances can be generated programmatically at scale without human annotation or LLM synthesis, allowing for virtually unlimited training data. ❷ Verifiability. SAT is a well-established NP-complete problem in theoretical computer science [SAT]. The correctness of a solution can be easily verified in linear time. But solving SAT problems requires complex reasoning. ❸ Controllable Difficulty. The difficulty of SAT instances can be precisely adjusted (e.g., number of variables, clauses), making it suitable for curriculum learning. What’s more, SAT serves as a universal substrate for limited forms of logical reasoning, as many problems in propositional logic, finite-domain first-order logic, and modal logic can be systematically reduced to SAT [logic1, logic2, logic3].
@@
We apply SATURN to DeepSeek-R1-Distill-Qwen-1.5B and 7B [deepseek_r1], obtaining SATURN-1.5B and SATURN-7B. Experiments show that SATURN effectively enhances LLMs’ reasoning capability in generalizable scenarios:
@@
  • The reasoning capability learned from SATURN transfers well to math and programming tasks, bringing average improvements of +4.9 and +1.8 on popular benchmarks such as AIME [aime], AMC [amc], MATH-500 [math500], GPQA Diamond [GPQA], and LiveCodebench [lcb] for two SATURN models.
@@ 2.1 SATURN Learning Loop Framework
Step 2: LLMs Training Loop. For the current difficulty, SAT_Construction generates a set of training instances that are different from the validation set. These samples are then used to train LLMs with GRPO. The reward function encourages outputs that are both logically correct and properly formatted. The training loop proceeds until $\texttt{pass@1}>\epsilon$ on the validation set. After that, the process backs to Step 1 to reassess and potentially advance the curriculum.
@@ 2.2 SAT Instances Construction
In this subsection, we formalize the construction of SAT instances. A SAT problem determines whether a propositional formula can be satisfied by a Boolean truth assignment. Formally, we define a $(n,k,l)$-SAT instance in conjunctive normal form (CNF) as:
$\left\{\begin{aligned} \Phi=\ &\left(x_{a_{1,1}}\lor\neg x_{a_{1,2}}\lor\dots\lor x_{a_{1,n}}\right)\land\cdots\land\left(x_{a_{l,1}}\lor\dots\lor\neg x_{a_{l,n}}\right)\\ \text{where }&a_{i,j}\in\{1,\dots,k\},\quad i\in[1,l]_{\mathbb{Z}},j\in[1,n]_{\mathbb{Z}}\end{aligned}\right$. (1)
where each clause contains exactly $n$ variables (literals), each being either $x_{i}$ or its negation $\neg x_{i}$, $k$ is the total number of variables, and $l$ is the total number of clauses. Based on the definition, we design a SAT instance constructor, SAT_Construction(n,k,l,m)(n,k,l,m), which uniformly samples $m$ SAT instances from the space of $(n,k,l)$-SAT. By adjusting the parameters $(n,k,l,m)$, SAT_Construction enables the scalable and controllable construction of SAT instances. The design details of the constructor algorithm are provided in Appendix B. All generated SAT instances are guaranteed to be satisfiable.
@@ 2.3 Estimation of Task Difficulty
As a canonical NP-complete problem [SAT], SAT admits a polynomial-time reduction from any other NP problem [reduction]. SAT exhibits a known phase transition phenomenon: when the clause-to-variable ratio $\alpha_{c}=l/k$ approaches a critical threshold (typically near $4.26$ for 3-SAT), the probability of satisfiability drops sharply, and problem difficulty peaks. This phenomenon probably stems from replica symmetry breaking (RSB) [SAT-RSB]: near $\alpha_{c}$, the solution space fractures into disconnected clusters separated by energy barriers. Beyond $\alpha_{c}$, the space collapses, reducing complexity.
However, RSB theory is designed for heuristic SAT solvers. For humans or LLMs solving SAT problems through logical steps such as trial, verification, and reasoning, such solver-like phase transitions are hardly observable in human-like thinking processes. While any $n$-SAT ($n>3$) can be reduced to 3-SAT [reduction], they differ significantly for LLMs in terms of solution space size and token length.
Prior work [sat_iclr_1] on SAT tasks for LLMs typically categorized difficulty based on phase transition points. To systematically estimate task difficulty, we define an analytical estimator of the expected solution space size. Given a $(n,k,l)$ SAT instance, its difficulty for LLMs can be approximately estimated by:
@@
Eq. (2) provides a more controllable, fine-grained estimation of SAT task difficulty. The detailed derivation is provided in Appendix C. To further validate Eq. (2), we evaluate LLMs’ performance on SAT instances with varying difficulty levels. As shown in Figure 3, each point represents a LLM’s average pass@3 on the same estimated difficulty instances. Pass@3 generally decreases as $D(n,k,l)$ increases, suggesting that our estimation aligns with the solvability trends observed in practical LLMs. Stronger LLMs maintain higher pass@3, while weaker LLMs exhibit lower scores overall. The validity of the estimation in Eq. (2) is further confirmed by ablation experiments, as detailed in Appendix C.
@@ 2.4 Reinforcement Learning with GRPO
In this subsection, we introduce the single-stage RL training for given $(n,k,l)$-difficulty tasks. RL can further improve LLMs’ generalization by directly optimizing policy gradients over diverse reasoning trajectories [deepseek_r1]. Given the SAT tasks, we then train LLMs using the original sample-level GRPO to optimize the policy $\pi_{\theta}$ with KL divergence penalty. The GRPO objective function is defined as:
@@
$\text{where }r_{i,t}(\theta)$ $=\frac{\pi_{\theta}(o_{i,t}\mid q,o_{i,<t})}{\pi_{\theta_{\text{old}}}(o_{i,t}\mid q,o_{i,<t})},\qquad\hat{A}_{i,t}=\frac{r_{i}-\text{mean}(\mathbf{r})}{\text{std}(\mathbf{r})}$
@@ 3.2 Experimental Setup
❶Building upon SAT_Construction tool and difficulty estimation, we release SATURN-2.6k, a curated benchmark designed to evaluate LLMs’ reasoning capability across varying complexity. SATURN-2.6k consists of 1,500 training instances and 160 test instances that share the same estimated difficulty level. To assess performance under increasing task complexity, SATURN-2.6k further includes 1,000 test instances from 10 unseen harder difficulty levels, with 100 instances per level. These levels are selected based on our difficulty estimation $D(n,k,l)$, enabling a systematic analysis of how LLM performance changes as problem difficulty increases. Additionally, custom datasets of desired difficulty can be constructed using our open-sourced SAT_Construction tool. ❷ For math and programming tasks, following DeepSeek-AI [deepseek_r1], we use AIME 24/25 [aime], AMC 22/23 [amc], MATH-500 [math500], GPQA Diamond [GPQA], and LiveCodeBench v4_v5 subset [lcb].
@@
We conduct evaluations against several 1.5B and 7B parameter reasoning models as the baselines, which include DeepSeek-R1-Distill-Qwen-1.5B & 7B [deepseek_r1], Still-3-1.5B-Preview [still3], s1.1-1.5B & 7B [s1], z1-7B [z1], OpenThinker-7B [openthoughts], and DeepScaleR-1.5B-Preview [deepscaler2025]. In addition, we include a supervised fine-tuning (SFT)-only baseline trained on the Math training dataset [math500], which provides step-by-step problem reasoning trajectories. We randomly select the most difficult Level-5 1,000 problems from training set for one epoch of SFT, following the same training template as DeepSeek-R1-Distill-Qwen. With the same dataset size, our setup enables a fair comparison between SFT and RL on SAT tasks.
@@
Following DeepSeek-AI [deepseek_r1], we use pass@k as the evaluation metric. Pass@k assesses the probability that at least one correct solution is generated within $k$ attempts. For SAT problems, we evaluate $\texttt{pass@k}\in\{1,3,5,7,10\}$ and sample 12 times per problem. For math and programming benchmarks, we use pass@1, following a context length of 32,768 and temperature = 0.6. More evaluation hyperparameters are provided in Appendix E. All experiments are conducted on NVIDIA 8×A100 (40GB) GPUs. Specific prompts are detailed in Appendix F.
@@ 3.3 RQ1: SATURN Substantially Improves Performance on SAT Tasks
We evaluate the performance of SATURN-1.5B and SATURN-7B on SAT tasks using SATURN-2.6k test set. Specifically, the evaluation involves unseen SAT instances that were not included in the training data. The results, presented in Table 2, 1014, and detailed in Appendix G, demonstrate the performance of LLMs across different SAT difficulties.
@@ 3.4 RQ2: SATURN Demonstrates Strong Generalization to Math and Programming
❶SATURN shows strong generalization to math and programming tasks. On the AIME 24/25 benchmark, SATURN-1.5B outperforms z1-7B by 8.3 and s1.1-7B by 21.7. Similarly, SATURN-7B achieves a strong improvement on the Math500 dataset, increasing from 93.2 to 95.0. On LiveCodeBench, it improves from 35.4 to 37.7. On average, SATURN-1.5B improves by +4.9, and SATURN-7B improves by +1.8 across these benchmarks. These results highlight that SATURN enhances the reasoning performance of LLMs across various math and programming tasks, demonstrating strong generalization of the learned reasoning capabilities from SAT.
❷SATURN outperforms SFT on broader benchmarks. Consistent with the observations in SFT Memorizes, RL Generalizes [SFT-RL], SFT improves performance on math-focused benchmarks (AIME, AMC, and Math500) that are similar to its supervised training domain. However, on LiveCodeBench, SFT drops from 16.4 to 14.6, exhibiting an alignment tax [alignment_tax], where specializing on a narrow domain compromises performance on other tasks. In contrast, SATURN improves performance across all benchmarks, with SATURN-1.5B reaching 17.4 on LiveCodeBench. Averaging across all benchmarks, SATURN-1.5B outperforms the SFT counterpart by 3.3, demonstrating that SATURN generalizes effectively.
@@ 3.5 RQ3: SATURN Serves as a Complement and Further Enhances LLM Reasoning
RQ3 studies the relationship between SATURN and existing RL tasks. Beyond the DeepSeek-R1-Distill-Qwen-7B, we introduce two additional models: Qwen2.5-7B-Instruct-1M [qwen2.5, qwen2.5-1m] following Logic-RL [Logic-RL] settings, and DeepScaleR-1.5B-Preview [deepscaler2025], which is further RL trained on 40k math and programming examples from DeepSeek-R1-Distill-Qwen-7B. We compare SATURN against several prior constructing RL task approaches, including Logic-RL [Logic-RL], SPGA [Self-playing-Adversarial], and ScaleQuest [question_synthesis], which represent strong baselines. Each approach is applied to different models for comparison. Results are summarized in Table 4.